# Classifying localizing subcategories of a Grothendieck category

### v1

@@ 5. Ziegler spectrum of a locally coherent category
Throughout this section $\mathcal{A}$ is a locally coherent category with a generating set.
[…]
$\mathcal{A}$ is locally monoform if and only $M$ is locally monoform for any $M\in\operatorname{fp-}\mathcal{A}$.
Proof.
The part "if" is clear. Conversely, given $M\in\mathcal{A}$, it contains a finitely generated subobject $N$. It is clear by section 4 that $N$ is the quotient of some finitely presented object $X$ in $\mathcal{A}$. Now the assumption implies that $N$ contains a monoform subobject. ∎
A category $\mathcal{A}$ is said to be locally uniform (or locally coirreducible by [Po, p. 330]) if every nonzero object in $\mathcal{A}$ has a uniform subobject. The following theorem establishes a characterization of locally monoform objects.
As the injective envelope of any monoform object is indecomposable, $\operatorname{ZASpec}\mathcal{A}$ can be considered as a subset of $\operatorname{Zg}\mathcal{A}$. Then we may identify $\operatorname{ZASpec}\mathcal{A}$ with a subset of $\operatorname{Zg}\mathcal{A}$ via the map $\alpha\mapsto E(\alpha)$. Furthermore, we have the following lemma.
[…]
@@
The injectivity of $\theta$ is clear. For every $M\in\operatorname{fp-}\mathcal{A}$, it follows from section 3 that $\theta^{-1}(\mathcal{O}(M))=\operatorname{ASupp}M$. Then $\theta^{-1}(\mathcal{O})$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any open subset $\mathcal{O}$ of $\operatorname{Zg}\mathcal{A}$. To prove the second claim, every finitely presented object $M$ has a maximal subobject $N$ so that $M/N$ is a monoform object in $\mathcal{A}$. If we consider $\alpha=\overline{M/N}$, it is clear that $\operatorname{Hom}(M,E(\alpha))\neq 0$. ∎
[…]
@@
Let $\mathcal{A}$ be locally monoform. The it is clear that $\mathcal{A}$ is locally uniform. Let $E$ be an indecomposable injective object in $\mathcal{A}$. By the assumption, the object $E$ contains a monoform subobject $H$ and so $E=E(\alpha)$, where $\alpha=\overline{H}$. This implies that $f$ is surjective. Therefore, it follows from section 5 that $f$ is a homeomorphism. Conversely, let $M$ be an object in $\mathcal{A}$. Since $\mathcal{A}$ is locally uniform, $M$ contains a uniform subobject $U$ so that $E(U)$ is an injective indecomposable. Since $f$ is surjective, there exists a monoform object $\alpha\in\operatorname{ASpec}\mathcal{A}$ such that $E(\alpha)=E(U)$. Then there exists a monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $E(H)=E(U)$ so that $H\cap U$ is a monoform subobject of $M$. ∎
[…]
The maps $\mathcal{O}(M)\rightarrow\operatorname{ASupp}M$ and $\operatorname{ASupp}M\rightarrow\mathcal{O}(M)$ establish a one to one correspondence between the lattices of open subsets of $\operatorname{Zg}\mathcal{A}$ and the lattices of open subsets of $\operatorname{ZASpec}\mathcal{A}$.
Proof.
This is clear. ∎
@@
[…]
Let $A=k[X_{n}(n\in\omega)]$ be a polynomial ring over a field $k$ in infinitely many commuting indeterminates. Then $A$ is clearly a commutative coherent ring. Let $I=(X_{n}^{n+1}:n\in\omega)$. Then $I$ is not prime as $X_{1}\notin I$ but $X_{1}^{2}\in I$. According to [P], the injective $A$-module $E(A/I)$ is indecomposable and it does not have the form $E(A/\mathfrak{p})$ for any prime ideal $\mathfrak{p}$ of $A$ by [P, 9.1]. In section 5, if $f$ is surjective, then there exists $\alpha\in\operatorname{ASpec}\operatorname{Mod}A$ such that $E(A/I)=E(\alpha)$. Hence there exists a cyclic monoform $A$-module $H$ such that $E(H)=E(A/I)$. But section 5 implies that $H=A/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$ of $A$ which is a contradiction. Therefore $\operatorname{Mod}A$ is not locally monoform by section 5. Furthermore, since $E(A/I)$ does not have the form $E(A/\mathfrak{p})$ for any prime ideal, by [GP, Theorem 1.4], there exists a prime ideal $\mathfrak{p}$ of $A$ such that $\varLambda(E(A/I))=\varLambda(E(A/\mathfrak{p}))$, where for any injective indecomposable $E$, the set $\varLambda(E)$ denotes the intersection of all open subsets of $\operatorname{Zg}\operatorname{Mod}A$ containing $E$. But this implies that $\operatorname{Zg}\operatorname{Mod}A$ is not $T_{0}$-space.
In the rest of this section we provide a new classification of localizing subcategories of finite type of $\mathcal{A}$. For a locally coherent category $\mathcal{A}$, Krause [Kr] has constructed a topology on $\operatorname{Sp}\mathcal{A}$ in which for a subset $\mathcal{U}$ of $\operatorname{Sp}\mathcal{A}$, the closure of $\mathcal{U}$ is defined as $\overline{\mathcal{U}}=\left\langle{}^{\bot}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}$. The subsets $\mathcal{U}$ of $\operatorname{Sp}\mathcal{A}$ satisfying $\mathcal{U}=\overline{\mathcal{U}}$ form the closed subsets of $\operatorname{Sp}\mathcal{A}.$ We observe that $\operatorname{Zg}\mathcal{A}$ and $\operatorname{Sp}\mathcal{A}$ have the same objects with relatively different topologies. The following proposition shows that the topologies of $\operatorname{Zg}\mathcal{A}$ and $\operatorname{Sp}\mathcal{A}$ are identical.
[…]
The spectra $\operatorname{Zg}(\mathcal{A})$ and $\operatorname{Sp}\mathcal{A}$ have the same topologies.
@@
In the following theorem we classify localizing subcategories of finite type of $\mathcal{A}$ via their Ziegler supports which are open subsets of $\operatorname{Zg}\mathcal{A}$.
[…]
The map $\mathcal{X}\mapsto\operatorname{ZSupp}\mathcal{X}$ provides a bijective correspondence between the lattice of localizing subcategories of finite type of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Zg}\mathcal{A}$. The inverse map is $\mathcal{U}\mapsto\operatorname{ZSupp}^{-1}\mathcal{U}$.
@@
Given a localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, by [Kr, Lemma 2.3], we have $\mathcal{X}=\overset{\rightarrow}{\mathcal{S}}$, where $\mathcal{S}=\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$. Then $\operatorname{ZSupp}\mathcal{X}=\bigcup_{M\in\mathcal{S}}\operatorname{ZSupp}M=\bigcup_{M\in\mathcal{S}}\mathcal{O}(M)$ that is an open subset of $\operatorname{Zg}\mathcal{A}$. Given an open subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, it is clear by the definition that $\operatorname{ZSupp}^{-1}\mathcal{U}=^{\bot}\mathcal{U}^{c}$. Then [Kr, Corollary 4.3] and section 5 imply that $\operatorname{ZSupp}^{-1}\mathcal{U}$ is a localizing subcategory of finite type of $\mathcal{A}$. To prove $\mathcal{X}=\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$, clearly $\mathcal{X}\subset\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$. For the other side, by the previous argument, $\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$ is a localizing subcategory of finite type of $\mathcal{A}$. Thus, for every $M\in\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$, we have $M=\underset{\rightarrow}{\rm lim}M_{i}$ where each $M_{i}$ belongs to $\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$ by [Kr, Lemma 2.3]. Then $\operatorname{ZSupp}M_{i}\subset\operatorname{ZSupp}\mathcal{X}$ for each $i$. Fixing $i$, since $\operatorname{ZSupp}M_{i}$ is a quasi-compact open subset of $\operatorname{Zg}\mathcal{A}$ by [Kr, Corollary 4.6], there exists $N\in\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$ such that $\operatorname{ZSupp}M_{i}\subset\operatorname{ZSupp}N$ and hence it follows from [H, Corollary 3.12] that $M_{i}\in\surd{N}$, where $\surd{N}$ is the smallest Serre subcategory of $\operatorname{fp-}\mathcal{A}$ containing $N$. Clearly $\surd{N}\subset\mathcal{X}$ and hence $M_{i}\in\mathcal{X}$. Finally, this forces that $M\in\mathcal{X}$ as $M=\underset{\rightarrow}{\rm lim}M_{i}$. It is clear that $\operatorname{ZSupp}(\operatorname{ZSupp}^{-1}\mathcal{U})\subset\mathcal{U}$. On the other hand, for every $I\in\mathcal{U}$, there exists a finitely presented object $M$ such that $I\in\mathcal{O}(M)=\operatorname{ZSupp}M\subset\mathcal{U}$. This implies that $M\in\operatorname{ZSupp}^{-1}\mathcal{U}$ and so the previous argument forces that $\operatorname{ZSupp}M\subset\operatorname{ZSupp}(\operatorname{ZSupp}^{-1}\mathcal{U})$. Therefore $I\in\operatorname{ZSupp}(\operatorname{ZSupp}^{-1}\mathcal{U})$; and hence $\mathcal{U}=\operatorname{ZSupp}(\operatorname{ZSupp}^{-1}\mathcal{U})$. ∎
[…]
The map $\mathcal{X}\mapsto\operatorname{ZSupp}\mathcal{X}$ provides a bijective correspondence between the lattice of Serre subcategories of $\operatorname{fp-}\mathcal{A}$ and the lattice of open subsets of $\operatorname{Zg}\mathcal{A}$. The inverse map is $\mathcal{U}\mapsto\operatorname{ZSupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}$.
@@
Given a Serre subcategory $\mathcal{X}$ of $\operatorname{fp-}\mathcal{A}$, we prove that $\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}=\mathcal{X}$. The inclusion $\mathcal{X}\subset\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$ is clear. For the other side, if $M\in\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})\cap\operatorname{fp-}\mathcal{A}$, then $\operatorname{ZSupp}M\subset\operatorname{ZSupp}\mathcal{X}\subset\operatorname{ZSupp}\overset{\longrightarrow}{\mathcal{X}}$. Thus section 5 implies that $M\in\overset{\longrightarrow}{\mathcal{X}}$ as $\mathcal{X}$ is a localizing subcategory of finite type of $\mathcal{A}$ by [Kr, Theorem 2.8]. Thus $M$ is a direct summand of some object $X\in\mathcal{X}$ so that $M\in\mathcal{X}$. Given an open subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, by the same proof as used in section 5, we deduce that $\operatorname{ZSupp}(\operatorname{ZSupp}^{-1}\mathcal{U}\cap\operatorname{fp-}\mathcal{A})=\mathcal{U}$. ∎
The above theorem provides a characterization of localizing subcategories of finite type of $\mathcal{A}$.
[…]
Let $\mathcal{X}$ be a localizing subcategory of $\mathcal{A}$. Then $\mathcal{X}$ is of finite type if and only if $\operatorname{ZSupp}\mathcal{X}$ is an open subset of $\operatorname{Zg}\mathcal{A}$.
@@
The part "only if " is clear. Conversely, if $\operatorname{ZSupp}\mathcal{X}$ is an open subset of $\operatorname{Zg}\mathcal{A}$, then a similar proof of section 5 shows that $\mathcal{X}=\operatorname{ZSupp}^{-1}(\operatorname{ZSupp}\mathcal{X})$. Consequently, $\mathcal{X}$ is of finite type of $\mathcal{A}$ by section 5. ∎
6. In the case of commutative (coherent) rings
In this section, we assume that $A$ is a commutative ring and we denote the category of $A$-modules by $\operatorname{Mod}A$. Also, we denote $\operatorname{ASpec}\operatorname{Mod}A$, $\operatorname{ZASpec}\operatorname{Mod}A$, $\operatorname{Zg}\operatorname{Mod}A$ and $\operatorname{Sp}\operatorname{Mod}A$, by $\operatorname{ASpec}A$, $\operatorname{ZASpec}A$, $\operatorname{Zg}A$ and $\operatorname{Sp}A$, respectively. The ring $A$ is said to be semi-noetherian (respectively, locally monoform) if $\operatorname{Mod}A$ is a semi-noetherain (respectively, locally monoform) category. We have the following result.
[…]
If $A$ is a semi-noetherian ring, then it is locally monoform. Moreover, if $A/\mathfrak{p}$ has Krull-Gabriel dimension for any $\mathfrak{p}\in\operatorname{Spec}A$, then the converse holds.
Proof.
See section 2 and [Po, p. 352, Theorem 5.16] ∎
[…]
The set =$\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is an ideal of $\mathcal{A}\}$ forms a basis for $\operatorname{ASpec}A$.
Proof.
Let $\mathfrak{a}$ and $\mathfrak{b}$ be ideals of $A$ and let $\alpha\in\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}$. By section 5, there exists a prime ideal $\mathfrak{p}$ such that $\overline{A/\mathfrak{p}}=\alpha$. It follows from section 3 that $\operatorname{Hom}_{A}(A/\mathfrak{a},E(A/\mathfrak{p}))\neq 0$ and $\operatorname{Hom}_{A}(A/\mathfrak{b},E(A/\mathfrak{p}))\neq 0$. This implies that $\mathfrak{p}\in V(\mathfrak{a})\cap V(\mathfrak{b})=V(\mathfrak{a}+\mathfrak{b})$; and hence we have $\alpha\in\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. Therefore $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}\subseteq\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$ and a similar argument implies that $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}=\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. On the other hand, assume $M$ is finitely generated, say $M=<x_{1},\dots,x_{n}>$. Set $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}$ is cyclic for each $i$. Thus there exist ideals $\mathfrak{a}_{i}$ of $A$ such that $M_{i}/M_{i-1}\cong A/\mathfrak{a}_{i}$ for all $1\leq i\leq n$. It is clear that $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\operatorname{ASupp}A/\mathfrak{a}_{i}$. Now, if $M$ is an arbitrary $A$-module, then $M$ is the direct union of its finitely generated submodules. Then, in view of the previous argument, there exists a set $\varLambda$ such that $\operatorname{ASupp}M=\bigcup_{\lambda\in\varLambda}\operatorname{ASupp}A/\mathfrak{a}_{\lambda}$. ∎
[…]

### v2

@@ 4. Ziegler spectrum of a locally coherent category
Throughout this section $\mathcal{A}$ is a locally coherent. The collection of the isomorphism classes of indecomposable injective objects in $\mathcal{A}$, denoted by $\operatorname{Zg}\mathcal{A}$, forms a set. This is because every indecomposable injective object is the injective envelope of some quotient of an element of the generating set of $\mathcal{A}$.
Definition 4.1.
The Ziegler support $\operatorname{\mathcal{O}}(M)$ of an object $M$ in $\mathcal{A}$ is a subset of $\operatorname{Zg}\mathcal{A}$, that is
$\operatorname{\mathcal{O}}(M)=\{I\in\operatorname{Zg}\mathcal{A}|\hskip 2.84544pt\operatorname{Hom}(M,I)\neq 0\}.$
For any exact sequence $0\rightarrow N\rightarrow M\rightarrow K\rightarrow 0$ of objects in $\mathcal{A}$, we have
$\operatorname{\mathcal{O}}(M)=\operatorname{\mathcal{O}}(N)\cup\operatorname{\mathcal{O}}(K).$
For any subcategory $\mathcal{X}$ of $\mathcal{A}$, we define $\operatorname{\mathcal{O}}(\mathcal{X})=\bigcup_{M\in\mathcal{X}}\operatorname{\mathcal{O}}(M)$ and for any subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, we define $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})=\{M\in\mathcal{A}|\hskip 2.84544pt\operatorname{\mathcal{O}}(M)\subset\mathcal{U}\}$.
Herzog [H] proved that the collection $\{\mathcal{O}(M)|\hskip 2.84544ptM\in\operatorname{fp-}\mathcal{A}\}$ satisfies the axioms for a basis of open subsets defining a topology on $\operatorname{Zg}\mathcal{A}$. The resulting topological space is called the Ziegler spectrum of $\mathcal{A}$.
For $\alpha\in\operatorname{ASpec}\mathcal{A}$ and monoform objects $H_{1}$ and $H_{2}$ in $\mathcal{A}$ satisfying $\alpha=\overline{H_{1}}=\overline{H_{2}}$, we have $E(H_{1})=E(H_{2})$. The isomorphism class of all such $E(H)$ such that $\alpha=\overline{H}$ is denoted by $E(\alpha)$. We observe that $E(\alpha)$ is an indecomposable injective object in $\mathcal{A}$. Because if $E(\alpha)=E(H)$ for some monoform object $H$ in $\mathcal{A}$ and $E(\alpha)=E_{1}\oplus E_{2}$, then $E_{1}\cap H$ and $E_{2}\cap H$ are nonzero monoform subobjecs of $H$. Thus $E_{1}\cap H\cap E_{2}\cap H$ is nonzero which is a contradiction. Then we may identify $\operatorname{ZASpec}\mathcal{A}$ with a subset of $\operatorname{Zg}\mathcal{A}$ via the map $\alpha\mapsto E(\alpha)$. Furthermore, we have the following lemma.
Lemma 4.2.
@@
The injectivity of $\theta$ is clear. For every $M\in\operatorname{fp-}\mathcal{A}$, it follows from section˜3 that $\theta^{-1}(\mathcal{O}(M))=\operatorname{ASupp}M$. Then $\theta^{-1}(\mathcal{O})$ is an open subset of $\operatorname{ZASpec}\mathcal{A}$ for any open subset $\mathcal{O}$ of $\operatorname{Zg}\mathcal{A}$. To prove the second claim, every finitely presented object $M$ has a maximal subobject $N$ so that $M/N$ is a monoform object in $\mathcal{A}$. If we consider $\alpha=\overline{M/N}$, it is clear that $\operatorname{Hom}(M,E(\alpha))\neq 0$. ∎
A category $\mathcal{A}$ is said to be locally uniform (or locally coirreducible by [Po, p. 330]) if every nonzero object in $\mathcal{A}$ has a uniform subobject. The following proposition establishes a characterization of locally monoform objects.
Proposition 4.3.
@@
Let $\mathcal{A}$ be locally monoform. The it is clear that $\mathcal{A}$ is locally uniform. Let $E$ be an indecomposable injective object in $\mathcal{A}$. By the assumption, the object $E$ contains a monoform subobject $H$ and so $E=E(\alpha)$, where $\alpha=\overline{H}$. This implies that $f$ is surjective. Therefore, it follows from section˜4 that $f$ is a homeomorphism. Conversely, let $M$ be an object in $\mathcal{A}$. Since $\mathcal{A}$ is locally uniform, $M$ contains a uniform subobject $U$ so that $E(U)$ is an injective indecomposable. Since $f$ is surjective, there exists a monoform object $\alpha\in\operatorname{ASpec}\mathcal{A}$ such that $E(\alpha)=E(U)$. Then there exists a monoform object $H$ in $\mathcal{A}$ such that $\alpha=\overline{H}$ and $E(H)=E(U)$ so that $H\cap U$ is a monoform subobject of $M$. ∎
@@
Example 4.5.
Let $A=k[X_{n}(n\in\omega)]$ be a polynomial ring over a field $k$ in infinitely many commuting indeterminates. Then $A$ is clearly a commutative coherent ring. Let $I=(X_{n}^{n+1}:n\in\omega)$. Then $I$ is not prime as $X_{1}\notin I$ but $X_{1}^{2}\in I$. According to [P], the injective $A$-module $E(A/I)$ is indecomposable and it does not have the form $E(A/\mathfrak{p})$ for any prime ideal $\mathfrak{p}$ of $A$ by [P, 9.1]. In section˜4, if $f$ is surjective, then there exists $\alpha\in\operatorname{ASpec}\operatorname{Mod}A$ such that $E(A/I)=E(\alpha)$. Hence there exists a cyclic monoform $A$-module $H$ such that $E(H)=E(A/I)$. But section˜4 implies that $H=A/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$ of $A$ which is a contradiction. Therefore $\operatorname{Mod}A$ is not locally monoform by section˜4. Furthermore, since $E(A/I)$ does not have the form $E(A/\mathfrak{p})$ for any prime ideal, by [GP, Theorem 1.4], there exists a prime ideal $\mathfrak{p}$ of $A$ such that $\varLambda(E(A/I))=\varLambda(E(A/\mathfrak{p}))$, where for any injective indecomposable $E$, the set $\varLambda(E)$ denotes the intersection of all open subsets of $\operatorname{Zg}\operatorname{Mod}A$ containing $E$. But this implies that $\operatorname{Zg}\operatorname{Mod}A$ is not $T_{0}$-space.
The following proposition provide a sufficient condition under which a locally coherent category is semi-noetherian (and so locally monoform).
Proposition 4.6.
If $\mathcal{A}_{\tau}$ is a localizing subcategory of finite type of $\mathcal{A}$, then $\mathcal{A}$ is semi-noetherian.
Proof.
If $\mathcal{A}$ is not semi noetherian, then $\mathcal{A}_{\tau}$ is a proper subcategory of $\mathcal{A}$. It follows from [H, Theorem 2.16] that $\mathcal{A}/\mathcal{A}_{\tau}$ is locally coherent so that it contains a simple object $S$. Hence there exists an object $X\in\mathcal{A}_{\tau+1}\setminus\mathcal{A}_{\tau}$ such that $F_{\tau}(X)=S$ which is a contradiction. ∎
Krause [Kr] has constructed a topology on $\operatorname{Zg}\mathcal{A}$ in which for a subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, the closure of $\mathcal{U}$ is defined as $\overline{\mathcal{U}}=\left\langle{}^{\bot}\mathcal{U}\cap\operatorname{fp-}\mathcal{A}\right\rangle^{\bot}$. The subsets $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$ satisfying $\mathcal{U}=\overline{\mathcal{U}}$ form the closed subsets of $\operatorname{Zg}\mathcal{A}.$ The following proposition shows that the Ziegler topology and the topology defined by Krause are identical.
Lemma 4.7.
The Ziegler topology and Krause topology on $\operatorname{Zg}(\mathcal{A})$ are the same.
@@
The localizing subcategories of finite type of $\mathcal{A}$ can be classifies via their Ziegler supports which are open subsets of $\operatorname{Zg}\mathcal{A}$.
Proposition 4.8.
The map $\mathcal{X}\mapsto\operatorname{\mathcal{O}}(\mathcal{X})$ provides a bijective correspondence between the lattice of localizing subcategories of finite type of $\mathcal{A}$ and the lattice of open subsets of $\operatorname{Zg}\mathcal{A}$. The inverse map is $\mathcal{U}\mapsto\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$.
@@
Given a localizing subcategory $\mathcal{X}$ of finite type of $\mathcal{A}$, by [Kr, Lemma 2.3], we have $\mathcal{X}=\overset{\rightarrow}{\mathcal{S}}$, where $\mathcal{S}=\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$. Then $\operatorname{\mathcal{O}}(\mathcal{X})=\bigcup_{M\in\mathcal{S}}\operatorname{\mathcal{O}}(M)$ that is an open subset of $\operatorname{Zg}\mathcal{A}$. Given an open subset $\mathcal{U}$ of $\operatorname{Zg}\mathcal{A}$, it is clear by the definition that $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})=^{\bot}\mathcal{U}^{c}$. Then [Kr, Corollary 4.3] and section˜4 imply that $\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$ is a localizing subcategory of finite type of $\mathcal{A}$. To prove $\mathcal{X}=\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$, clearly $\mathcal{X}\subset\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$. For the other side, by the previous argument, $\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$ is a localizing subcategory of finite type of $\mathcal{A}$. Thus, for every $M\in\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))$, we have $M=\underset{\rightarrow}{\rm lim}M_{i}$ where each $M_{i}$ belongs to $\operatorname{\mathcal{O}}^{-1}(\operatorname{\mathcal{O}}(\mathcal{X}))\cap\operatorname{fp-}\mathcal{A}$ by [Kr, Lemma 2.3]. Then $\operatorname{\mathcal{O}}(M_{i})\subset\operatorname{\mathcal{O}}(\mathcal{X})$ for each $i$. Fixing $i$, since $\operatorname{\mathcal{O}}(M_{i})$ is a quasi-compact open subset of $\operatorname{Zg}\mathcal{A}$ by [Kr, Corollary 4.6], there exists $N\in\mathcal{X}\cap\operatorname{fp-}\mathcal{A}$ such that $\operatorname{\mathcal{O}}(M_{i})\subset\operatorname{\mathcal{O}}(N)$ and hence it follows from [H, Corollary 3.12] that $M_{i}\in\surd{N}$, where $\surd{N}$ is the smallest Serre subcategory of $\operatorname{fp-}\mathcal{A}$ containing $N$. Clearly $\surd{N}\subset\mathcal{X}$ and hence $M_{i}\in\mathcal{X}$. Finally, this forces that $M\in\mathcal{X}$ as $M=\underset{\rightarrow}{\rm lim}M_{i}$. It is clear that $\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))\subset\mathcal{U}$. On the other hand, for every $I\in\mathcal{U}$, there exists a finitely presented object $M$ such that $I\in\mathcal{O}(M)\subset\mathcal{U}$. This implies that $M\in\operatorname{\mathcal{O}}^{-1}(\mathcal{U})$ and so the previous argument forces that $\operatorname{\mathcal{O}}M\subset\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$. Therefore $I\in\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$; and hence $\mathcal{U}=\operatorname{\mathcal{O}}(\operatorname{\mathcal{O}}^{-1}(\mathcal{U}))$. ∎
5. In the case of commutative coherent rings
In this section, we assume that $A$ is a commutative coherent ring and we denote the category of $A$-modules by $\operatorname{Mod}A$. Also, we denote $\operatorname{ZASpec}\operatorname{Mod}A$, $\operatorname{Zg}\operatorname{Mod}A$ and $\operatorname{fp-}\operatorname{Mod}A$ by $\operatorname{ZASpec}A$, $\operatorname{Zg}A$ and $\operatorname{fp-}A$, respectively.
Lemma 5.1.
$\operatorname{ZASpec}\mathcal{A}$ is a retract subspace of $\operatorname{Zg}\mathcal{A}$.
@@ 5. In the case of commutative coherent rings
We define $\theta:\operatorname{Zg}A\rightarrow\operatorname{ZASpec}A$ given by $E\rightarrow E(A/{\mathfrak{p}_{E}})$. If $E(\alpha)\in\operatorname{ZASpec}\mathcal{A}$, then there exists a prime ideal $\mathfrak{p}=\mathfrak{p}_{E(\alpha)}$ of $A$ such that $\alpha=\overline{A/\mathfrak{p}}$ by section˜4. Hence $E(\alpha)=E(A/\mathfrak{p})$ so that $\theta(E(\alpha))=E(\alpha)$. For every finitely presented $A$-module $M$, we show that $\theta^{-1}(\operatorname{ASupp}M)=\mathcal{O}(M)$. If $E\in\theta^{-1}(\operatorname{ASupp}M)$, we have $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ so that $E(A/\mathfrak{p}_{E})\in\mathcal{O}(M)$ by section˜3. It now follows from [GP, Theorem 1.4] that $E\in\mathcal{O}(M)$. Conversely if $E\in\mathcal{O}(M)$, then $\theta(E)=E(A/\mathfrak{p}_{E})\in\operatorname{ASupp}M$ by [GP, Theorem 1.4]. Therefore $E\in\theta^{-1}(\operatorname{ASupp}M)$. ∎
Lemma 5.2.
The set $\mathcal{F}\mathcal{I}=\{\operatorname{ASupp}A/\mathfrak{a}|\hskip 2.84544pt\mathfrak{a}$ is a finitely generated ideal of A}A\} forms a basis for the topology $\operatorname{ZASpec}A$.
@@
We observe that $\operatorname{ASpec}A=\operatorname{ASupp}A$. For any finitely generated ideals $\mathfrak{a}$ and $\mathfrak{b}$ of $A$ and $\alpha\in\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}$. By section˜4, there exists a prime ideal $\mathfrak{p}$ such that $\overline{A/\mathfrak{p}}=\alpha$. It follows from section˜3 that $\operatorname{Hom}_{A}(A/\mathfrak{a},E(A/\mathfrak{p}))\neq 0$ and $\operatorname{Hom}_{A}(A/\mathfrak{b},E(A/\mathfrak{p}))\neq 0$. This implies that $\mathfrak{p}\in V(\mathfrak{a})\cap V(\mathfrak{b})=V(\mathfrak{a}+\mathfrak{b})$; and hence we have $\alpha\in\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. Therefore $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}\subseteq\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$ and an easy argument implies that $\operatorname{ASupp}A/\mathfrak{a}\cap\operatorname{ASupp}A/\mathfrak{b}=\operatorname{ASupp}A/(\mathfrak{a}+\mathfrak{b})$. On the other hand, assume that $M$ is a finitely presented $A$-module. Then $M$ is finitely generated, say $M=<x_{1},\dots,x_{n}>$. Set $M_{0}=0$ and $M_{k}=<x_{1},\dots x_{k}>$ for $1\leq k\leq n$. Then $M_{i}/M_{i-1}$ is cyclic and finitely presented for each $i$. Thus there exist finitely generated ideals $\mathfrak{a}_{i}$ such that $M_{i}/M_{i-1}\cong A/\mathfrak{a}_{i}$ for all $1\leq i\leq n$. It is clear that $\operatorname{ASupp}M=\bigcup_{i=1}^{n}\operatorname{ASupp}A/\mathfrak{a}_{i}$. ∎
Proposition 5.3.